Mr. Jackson owns a dairy farm that is bordered by a river on one side. He wants to enclose a rectangular pasture that is bordered on one side by a river and on the other three sides by 600 feet of fencing. What must the dimensions of the pasture be in order for Mr. Jackson to get the greates amount of grazing area, and what is the greatest area for grazing?
A=l*w
A=x(600-2x)

Answer by Alan3354(69443)

Mr. Jackson owns a dairy farm that is bordered by a river on one side. He wants to enclose a rectangular pasture that is bordered on one side by a river and on the other three sides by 600 feet of fencing. What must the dimensions of the pasture be in order for Mr. Jackson to get the greates amount of grazing area, and what is the greatest area for grazing?
A=l*w
A=x(600-2x)
A = 600x - 2x^2
Set the 1st derivative (dA/dx) equal to zero
600 - 4x = 0
x = 150 feet
It's 150 by 300 feet for max area.
Area = 45,000 sq ft